# Job-SDF: A Multi-Granularity Dataset for Job Skill Demand Forecasting and Benchmarking

### v2

@@ Appendix C Data Structure and Components
Our dataset comprises five components for each granularity level: job skill demand sequences, job skill demand proportion sequences, ID mapping index, the indexes of skills with structural breaks, and skill co-occurrence graph. Each component is structured as follows: (1) Job Skill Demand Sequences: These are presented in tabular files, where each row represents a specific skill, and each column corresponds to a different time slice (month). Each cell within the table contains a numerical value that reflects the demand for the respective skill during that month. (ii) Job Skill Demand Proportion Sequences: This component is also formatted in tabular files similar to the skill demand sequences. However, each cell in these tables displays a value between 0 and 1, representing the proportion of demand defined in Eq 6. This provides a normalized view of skill demand across different granularities. (iii) ID mapping index: In the dataset, various elements such as regions, occupations, companies, and skills are represented using unique identifiers (IDs) for the convenience of experimentation and analysis. An index table is provided that maps each ID to the actual names of regions, occupations, and skills, facilitating clear and effective data interpretation. The names of companies, however, are withheld due to potential privacy concerns. This selective anonymization ensures that while company-related data remains confidential, researchers can still access necessary details about other variables. During the review period, we open part of id-to-entity mappings to illustrate its use and structure. (iv) Indexes of skills with structural breaks: In the provided dataset, data concerning skills that have experienced structural breaks are organized in JSON format. Each granularity level is represented by a separate JSON file, which contains a list of indexes. These indexes correspond to the skills that have undergone structural breaks and can be directly mapped to the skill indexes in the skill demand sequences. The purpose of supplying this data is to facilitate research on the demand trends of skills that have exhibited structural breaks, enabling a detailed analysis of their demand dynamics over time. (v) Skill Co-occurrence Graph: This data is provided as a set of triples (skill ID_1, skill ID_2, frequency of co-occurrence), forming a collection that outlines the co-occurrence relationships between skills. Each triple indicates how frequently two skills are mentioned or required together within the job advertisements in the training data, serving as a prior knowledge graph to enhance predictive modeling by capturing relationships between skills.

### v3

@@ Appendix C Data Structure and Components
Our dataset comprises five components for each granularity level: job skill demand sequences, job skill demand proportion sequences, ID mapping index, the indexes of skills with structural breaks, and skill co-occurrence graph. Each component is structured as follows: (1) Job Skill Demand Sequences: These are presented in tabular files, where each row represents a specific skill, and each column corresponds to a different time slice (month). Each cell within the table contains a numerical value that reflects the demand for the respective skill during that month. (ii) Job Skill Demand Proportion Sequences: This component is also formatted in tabular files similar to the skill demand sequences. However, each cell in these tables displays a value between 0 and 1, representing the proportion of demand defined in Eq 6. This provides a normalized view of skill demand across different granularities. (iii) ID mapping index: In the dataset, various elements such as regions, occupations, companies, and skills are represented using unique identifiers (IDs) for the convenience of experimentation and analysis. An index table is provided that maps each ID to the actual names of regions, occupations, and skills, facilitating clear and effective data interpretation. The names of companies, however, are withheld due to potential privacy concerns. The remaining mapping tables are not publicly available on GitHub. Researchers requiring access to these mapping tables, excluding company-related data, may contact the first author via email to submit a request. (iv) Indexes of skills with structural breaks: In the provided dataset, data concerning skills that have experienced structural breaks are organized in JSON format. Each granularity level is represented by a separate JSON file, which contains a list of indexes. These indexes correspond to the skills that have undergone structural breaks and can be directly mapped to the skill indexes in the skill demand sequences. The purpose of supplying this data is to facilitate research on the demand trends of skills that have exhibited structural breaks, enabling a detailed analysis of their demand dynamics over time. (v) Skill Co-occurrence Graph: This data is provided as a set of triples (skill ID_1, skill ID_2, frequency of co-occurrence), forming a collection that outlines the co-occurrence relationships between skills. Each triple indicates how frequently two skills are mentioned or required together within the job advertisements in the training data, serving as a prior knowledge graph to enhance predictive modeling by capturing relationships between skills.